It took nick two-thirds of an hour to complete his mathematics homework on Monday, one fourth of an hour on Tuesday.how many hours did it take him to complete his homework altogether?





What is asked?




What are the given facts?




What operation to be use?




What is the number of sentence?




Solution




Answer







1
Expert's answer
2021-11-10T18:11:03-0500

a)

Question:

how many hours did it take him to complete his homework altogether?


b)

Given:

It took Nick two-thirds "(\dfrac{2}{3})" of an hour to complete his mathematics homework on Monday, one fourth "(\dfrac{1}{4})" of an hour on Tuesday.


c)

We are to use addition.


d)

"\dfrac{2}{3}+\dfrac{1}{4}="

e)

Solution


"\dfrac{2}{3}+\dfrac{1}{4}=\dfrac{2(4)+1(3)}{12}=\dfrac{11}{12}"

f)

Answer

It took Nick eleven twelfths "(\dfrac{11}{12})" of an hour to complete his mathematics homework altogether.